OpenStudy (anonymous):

Simplify the sum. State any restrictions on the variables. 2/3 + 10x/x^2-9

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

Oh crap - this is really making it difficult to read....is ur Q: 2/3 + 10x/(x^2-9) right? Or there is no bracket and the 9 is outside the division?

OpenStudy (anonymous):

\[\frac{ x-2 }{ x+3 }+\frac{ 10x }{ x ^{2}- 9}\]

OpenStudy (superdavesuper):

unfortunately LaTex is not working :( Let me see if I can type it out with brackets: [(x-2)/(x+3)] + [10x/(x^2-9)] is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

sorry stuck in updating OS :( OK to simplify this, plz note x^2-a^2=(x+a)(x-a) so can u do the same for x^2-9?

OpenStudy (anonymous):

i don't think so.. i dont really get it

OpenStudy (superdavesuper):

(x^2-9) is in the bottom of ur Q right? so u need to group it and I gave the eqn to group it in my last post.... x^2-9=x^2-3^2=???